# Cellular Wireless Networks in the Upper Mid-Band

### 11 Dec 2023 v5

@@ III-E Effects of Interference
Fig. 6 shows the CDF of the signal-to-interference-noise ratio (SINR) and the corresponding data rate calculated using Eq. 1. As shown in Fig. 6a, SINRs are degraded at lower frequencies (6 and ) because the small number of antenna arrays produces wider beamwidths. In contrast, due to directional transmissions with narrow beamwidths, UEs at higher frequencies experience less interference. As a result, we observe in Fig. 6b that data rates at higher frequency are almost uniformly better than at lower frequency.
However, some cautions on this result should be taken: the simulation is only for outdoor users. Due to the high penetration loss, indoor users will still likely require lower frequencies. Moreover, our simulation does not have blockers other than buildings. Practical scenarios, particularly for urban users at the street level, will experience blocking by vehicles and people that can reduce coverage. Nevertheless, an interesting finding from this result is that dynamic frequency selection in the upper mid-band can enhance the coverage and capacity gains when interference levels vary considerably. For example, when systems at lower frequencies become interference-limited, users can be moved to higher frequencies, provided that the propagation is not blocked.
IV Interference with Incumbent Services
IV-A Commercial communication satellites

### 06 Mar 2024 v6

@@ 3.5 Effects of Interference
Fig. 7 shows the CDF of the signal-to-interference-noise ratio (SINR) and the corresponding data rate calculated using (1). As shown in Fig. 7a, SINRs are degraded at lower frequencies (6 and ) because the small number of antenna arrays produces wider beamwidths. In contrast, due to directional transmissions with narrow beamwidths, UEs at higher frequencies experience less interference.
As a result, we observe in Fig. 7b that data rates at higher frequency are almost uniformly better than those at lower frequency. Thus, the “best choice” always selects the high-frequency band. This result suggests yet another benefit of adaptive systems: in interference-limited scenarios, BSs or UEs can automatically choose bands where directionality can reduce interference.
Note that to clearly see the effect of interference, we have not added blocking in this simulation as we did earlier. If blocking were added, then the high frequency performance would degrade and not be uniformly better. The point is that adaptive systems can naturally select the optimal bands with multiple factors, including path loss, penetration, and directionality.
4 Interference with Incumbent Services
4.1 Commercial Communication Satellites